The length of a rectangle is four times the width. The area is 144 in.^2. Find the length and the width. Show your work.

Answer by checkley71(8403)

L*W=144
L=4W
4W*W=144
4W^2=144
W^2=144/4
W^2=36
W=SQRT36
W=6 ANSWER FOR THE WIDTH.
L=4*6
L=24 ANSWER FOR THE LENGTH.
PROOF
6*24=144
144=144